The initial equation is C4H8 -> 2 C2H4, (cyclobutane to ethylene), the initial temperature is 400k, the reaction is 1st order and the rate constant is .0736

If you start with .5M cyclobutane, how long will it take to produce .420M of the product, ethylene?
Answer #1

First order rate equation is as follows:

ln[A]=ln[A]o - kt

Here [A]o is initial concentration

[A] concentration after time t

k is rate constant

t is time

ln[0.420] = ln[0.5] - 0.736 s-1(t)

-0.8675=-0.693-0.736s-1 (t)

0.736s-1(t) = 0.1745

t = 0.1745 / 0.736 s-1

t = 0.237 s